# Empowering Snapshot Compressive Imaging: Spatial-Spectral State Space Model with Across-Scanning and Local Enhancement

### v1

@@ Abstract
Snapshot Compressive Imaging (SCI) relies on decoding algorithms such as CNN or Transformer to reconstruct the hyperspectral image (HSI) from its compressed measurement. Although existing CNN and Transformer-based methods have proven effective, CNNs are limited by their inadequate modeling of long-range dependencies, while Transformer ones face high computational costs due to quadratic complexity. Recent Mamba models have demonstrated superior performance over CNN and Transformer-based architectures in some visual tasks, but these models have not fully utilized the local similarities in both spatial and spectral dimensions. Moreover, the long-sequence modeling capability of SSM may offer an advantage in processing the numerous spectral bands for HSI reconstruction, which has not yet been explored. In this paper, we introduce a State Space Model with Across-Scanning and Local Enhancement, named ASLE-SSM, that employs a Spatial-Spectral SSM for global-local balanced context encoding and cross-channel interaction promoting. Specifically, we introduce local scanning in the spatial dimension to balance the global and local receptive fields, and then propose our across-scanning method based on spatial-spectral local cubes to leverage local similarities between adjacent spectral bands and pixels to guide the reconstruction process. These two scanning mechanisms extract the HSI’s local features while balancing the global perspective without any additional costs. Experimental results illustrate ASLE-SSM’s superiority over existing state-of-the-art methods, with an inference speed 2.4$\times$ faster than Transformer-based MST and saving 0.12 (M) of parameters, achieving the lowest computational cost and parameter count.
@@ 1 Introduction
Figure 1: PSNR-FLOPs comparisons of ASLE-SSM and previous End-to-End Transformer-based SOTA methods.
@@
Hyperspectral images (HSIs) contain a large number of spectral bands to store rich detail information and thus are widely used in various fields such as object tracking ot_1 ; ot_2 ; ot_3 , remote sensing chen2017denoising ; he2020non ; yuan2018hyperspectral , super-resolution dong2021model ; sr_2 ; sr_3 , and medical image processing mi_1 ; mi_2 ; mi_3 . Traditional imaging systems used for collecting HSIs involve scanning along the spatial or spectral dimensions but this imaging process is time-consuming and can not effectively capture dynamic objects. Recently, snapshot compressive imaging (SCI) systems have been developed sci_1 ; sci_2 as an alternative method for obtaining HSIs at video rate. SCI systems capture multiple image bands in a single snapshot, compressing the information of HSIs into a 2D measurement. sci_3 ; sci_4 ; sci_5 ; sci_6 . Among these systems, Coded Aperture Snapshot Spectral Imaging (CASSI) has emerged and become mainstream gap_tv ; icme_sci_1 . CASSI utilizes a physical mask to modulate signals from HSIs, shifting these modulated images to different spatial locations on the detector plane via a disperser mst ; icme_sci_2 . However, the performance of the CASSI system heavily depends on the efficiency of the reconstruction algorithm.
@@
In recent years, numerous reconstruction algorithms have been proposed, including model-based methods and deep learning-based methods. Traditional model-based methods typically attempt to model the inverse process of image reconstruction, using manually designed priors to derive the solution to an optimization problem. Although these methods are highly interpretable, they often require manual parameter tuning and tend to be slow. With the success of deep learning methods in various visual tasks, CNN and Transformer-based methods have been developed and applied to HSI reconstruction, achieving higher image reconstruction quality and speed than traditional algorithms. The CNN-based methods lambda ; tsa_net have an advantage in extracting local features from images, but they lack the ability to model long-range dependencies effectively. Transformer-based methods mst ; mst_pp ; dauhst ; cst realize a global receptive field, resulting in superior performance compared to CNNs but higher computational costs. Although some mechanisms can mitigate this issue, they do not fundamentally change the quadratic complexity problem caused by the attention mechanism in the Transformer architecture.
Figure 2: The flowchart of the proposed ASLE-SSM. The framework is arranged by adopting a U-shaped structure. (a) Spatial-Spectral SSM Block. This module mainly consists of a spatial SSM and a spectral SSM, depicted in the sub-flowchart (b) and (c), respectively. (d) The specific method of LE-SSM in the Spatial SSM Module. Scanning directions are divided into global and local branches. (e) The illustration of our Across-Scanning Mechanism. The scanning process is demonstrated by labeling the order of each data point.
@@
The limitation of Transformers in spectral SCI remains unresolved. However, in the realm of large language model and RGB image vision, an advanced solution for this problem exists, known as Mamba gu2023mamba . Through data-dependent selective SSMs with parallel scanning and hardware-wise strategies, Mamba establishes a global receptive field and achieves near-linear complexity. Mamba-based methods have been shown to outperform previous methods in various tasks due to their excellent long sequence modeling capabilities and lower computational costs mamba_survey ; vmamba2 ; mambair ; lightm-unet .
@@
Subsequently, many works have begun to explore the application and performance of the Mamba model in various visual tasks. Representative works, such as the visual state space model shi2024vmambair and vision Mamba vision_mamba , have demonstrated Mamba’s superiority for image processing, particularly in obtaining fewer parameters and faster speed. However, most Mamba-based models only conduct the scanning process along the spatial dimension, lacking cross-channel interaction, despite the importance of spectral information for HSI reconstruction. Recently, some Mamba-based methods have made progress in developing an SSM mechanism along the channel (spectral) dimension for certain visual tasks s2mamba ; huang2024spectral , achieving better results compared to scanning only along the spatial dimension. However, this scanning method does not consider the inherently local features of HSIs in both spatial and spectral dimensions.
@@
In this paper, we introduce ASLE-SSM, a spatial-spectral SSM with across-scanning and local enhancement, to improve HSI reconstruction in SCI. Our ASLE-SSM aims to strike a balance between global and local information utilization in HSIs. ASLE-SSM comprises two main components: spatial SSM and spectral-spatial cross SSM. The spatial SSM employs a combination of global and local scanning methods, enhancing row-by-row scanning with both global and local modeling. Meanwhile, the spectral-spatial cross SSM incorporates an innovative across-scanning mechanism, leveraging spatial-spectral local cubes to enhance the module’s ability to capture local details. Overall, our contributions are summarized as follows:
@@
  • We propose an across-scanning method that integrates local features from both spatial and spectral dimensions within a single scan. Additionally, we introduce a local scanning approach within the spatial SSM to further enhance local feature extraction.
  • Our network outperforms state-of-the-art (SOTA) methods on simulation and real datasets, delivering visually superior results. Importantly, it boasts the minimum parameter count, lowest computational cost, and fastest inference speed among current SOTA methods.
Figure 3: A Schematic diagram of the coded aperture snapshot spectral imaging (CASSI) system. The 3D HSI cube is first modulated by a physical mask and then the modulated data cube is sheared by a disperser. The 2D-coded measurement thus includes the information on the HSI data cube.
@@ 2 Preliminaries and Related Works
2.1 CASSI System
CASSI consists of a hardware encoder and a software decoder. The encoder compresses a 3D data cube into a 2D snapshot measurement and the decoder reconstructs HSIs from the measurement. An imaging process CASSI system is illustrated in Fig. 3. The 3D HSI data cube is denoted as $\mathbf{F}\in\mathbb{R}^{H\times W\times N_{\lambda}}$ where H, W, and $N_{\lambda}$ represent the height, width, and number of wavelengths of the 3D HSI, respectively. A physical mask $\mathbf{M}\in\mathbb{R}^{H\times W}$ is used to modulate the 3D HSI $\mathbf{F}$ as:
$\mathbf{F^{{}^{\prime}}}(:,:,n_{\lambda})=\mathbf{F^{{}^{\prime}}}(:,:,n_{\lambda})\odot\mathbf{M},$ (1)
where $\odot$ represents the element-wise product and $\mathbf{F^{{}^{\prime}}}\in\mathbb{R}^{H\times W\times N_{\lambda}}$ is the modulated 3D HSIs. After the modulated cube passes the disperser, $\mathbf{F^{{}^{\prime}}}$ is tilted and considered to be sheared along the y-axis. We use $\mathbf{F^{{}^{\prime\prime}}}\in\mathbb{R}^{H\times(W+d(N_{\lambda}-1))\times N_{\lambda}}$ to denote the tilted cube and assume $n_{\lambda}$ to be the reference wavelength. That is, image $\mathbf{F^{{}^{\prime}}}(:,:,n_{\lambda_{c}})$ is not sheared along the y-axis. Then we have:
$\mathbf{F^{{}^{\prime\prime}}}(u,v,n_{\lambda})=\mathbf{F^{{}^{\prime}}}(x,y+d(\lambda_{n}-\lambda_{c}),n_{\lambda}),$ (2)
where $(u,v)$ indicates the coordinate system on the detector plane, and $\lambda_{n}$ is the wavelength of channel $n_{\lambda}$. Here, $d(\lambda_{n}-\lambda_{c})$ signifies the spatial shifting for channel $n_{\lambda}$. Finally, the 2D compressed measurement $\mathbf{Y}\in\mathbb{R}^{H\times[W+d(N_{\lambda}-1)]}$ can be obtained by
@@ 2.1 CASSI System
2.2 State Space Models
Recently, researchers have made significant progress in state space models (SSMs), leading to the emergence of various methods and architectures based on SSMs. Among them, the structured state space model (S4) s4gu2021 becomes a pioneer work for its advantage in modeling long-range dependencies. S4 maps a 1D sequence $x(t)\in\mathbb{R}\rightarrow y(t)\in\mathbb{R}$ by an implicit latent state $h(t)\in\mathbb{R}^{N}$ and decodes it into an output sequence, which can be formulated as:
@@ 2.2 State Space Models
where $N$ is the state size, ${\rm\textbf{A}}\in\mathbb{R}^{N\times N}$, ${\rm\textbf{B}}\in\mathbb{R}^{N\times 1}$, ${\rm\textbf{C}}\in\mathbb{R}^{1\times N}$, and ${\rm\textbf{D}}\in\mathbb{R}$. After that, a discretization process is adopted to integrate Eq. (4) into the deep learning algorithm, using the zero-order hold (ZOH) rule, which is defined as follows:
@@
However, the constant dynamics in S4 prevent it from correctly selecting the context information. By introducing a selective mechanism, selective structured state space models (S6), also called Mamba mambagu2023mamba have been proposed to realize the data-dependent dynamics, thereby effectively filtering the context information and achieving a better performance in language and visual tasks.
2.3 Mamba in Vision
Motivated by Mamba’s excellent performance in language tasks, recent research has begun exploring its application in vision tasks and has achieved promising results. To address the unidirectional modeling and lack of positional awareness in the original Mamba model, Vision Mamba (Vim) vision_mamba incorporates the bidirectional SSMs for data-dependent global visual context modeling and position embeddings for location-aware visual recognition. VMamba liu2024vmamba introduces the Cross-Scan Module (CSM) to bridge the gap between 1-D array scanning and 2-D plain traversing. However, Vim and VMamba have not explored the role of SSM in the channel dimension.
More recently, some works have begun to apply the SSM mechanism to the channel dimension. CU-Mamba cumamba employs a Channel SSM component to preserve channel correlation features. S2Mamba s2mamba designs an extra scanning module on the spectral dimension for retrieving semantic properties in HSI data by a bi-directional interaction between each band. However, these methods directly conduct the scanning process along the spectral dimension and do not utilize the local similarity of HSIs in both spatial and spectral dimensions.
3 The Proposed ASLE-SSM Method
3.1 Architecture of ASLE-SSM

### v2

# Cross-Scan Mamba with Masked Training for Robust Spectral Imaging

@@ Abstract
Snapshot Compressive Imaging (SCI) enables fast spectral imaging but requires effective decoding algorithms for hyperspectral image (HSI) reconstruction from compressed measurements. Current CNN-based methods are limited in modeling long-range dependencies, while Transformer-based models face high computational complexity. Although recent Mamba models outperform CNNs and Transformers in RGB tasks concerning computational efficiency or accuracy, they are not specifically optimized to fully leverage the local spatial and spectral correlations inherent in HSIs. To address this, we propose the Cross-Scanning Mamba, named CS-Mamba, that employs a Spatial-Spectral SSM for global-local balanced context encoding and cross-channel interaction promotion. Besides, while current reconstruction algorithms perform increasingly well in simulation scenarios, they exhibit suboptimal performance on real data due to limited generalization capability. During the training process, the model may not capture the inherent features of the images but rather learn the parameters to mitigate specific noise and loss, which may lead to a decline in reconstruction quality when faced with real scenes. To overcome this challenge, we propose a masked training method to enhance the generalization ability of models. Experiment results show that our CS-Mamba achieves state-of-the-art performance and the masked training method can better reconstruct smooth features to improve the visual quality.
@@ 1 Introduction
Figure 1: PSNR-FLOPs comparisons of CS-Mamba and previous Deep Unfolding SOTA methods.
@@
Hyperspectral images (HSIs) contain a large number of spectral bands to store rich detail information and thus are widely used in various fields such as object tracking [25, 36, 40], remote sensing [9, 21, 49], super-resolution [12, 44, 20], and medical image processing [1, 30, 33]. Traditional imaging systems used for collecting HSIs involve scanning along the spatial or spectral dimensions but this imaging process is time-consuming and can not effectively capture dynamic objects. Recently, snapshot compressive imaging (SCI) systems have been developed [29, 41] as an alternative method for obtaining HSIs at video rate. SCI systems capture multiple image bands in a single snapshot, compressing the information of HSIs into a 2D measurement. [42, 51, 6, 15]. Among these systems, Coded Aperture Snapshot Spectral Imaging (CASSI) has emerged and become mainstream [50, 48]. CASSI utilizes a physical mask to modulate signals from HSIs, shifting these modulated images to different spatial locations on the detector plane via a disperser [3, 47]. However, the performance of the CASSI system heavily depends on the efficiency of the reconstruction algorithm.
@@
In recent years, numerous reconstruction algorithms have been proposed, including model-based methods and deep learning-based methods. Traditional model-based methods typically attempt to model the inverse process of image reconstruction, using manually designed priors to derive the solution to an optimization problem. Although these methods are highly interpretable, they often require manual parameter tuning and tend to be slow. With the success of deep learning methods in various visual tasks, CNN and Transformer-based methods have been developed and applied to HSI reconstruction, achieving higher image reconstruction quality and speed than traditional algorithms. The CNN-based methods [35, 32] have an advantage in extracting local features from images, but they lack the ability to model long-range dependencies effectively. Transformer-based methods [3, 4, 5, 2] realize a global receptive field, resulting in superior performance compared to CNNs but higher computational costs. Although some mechanisms can mitigate this issue, they do not fundamentally change the quadratic complexity problem caused by the attention mechanism in the Transformer architecture.
@@
The limitation of Transformers in spectral SCI remains unresolved. However, in the realm of large language model and RGB image vision, an advanced solution for this problem exists, known as Mamba [16]. Through data-dependent selective SSMs with parallel scanning and hardware-wise strategies, Mamba establishes a global receptive field and achieves near-linear complexity. Mamba-based methods have been shown to outperform previous methods in various tasks due to their excellent long sequence modeling capabilities and lower computational costs [45, 54, 19, 27].
@@
Subsequently, many works have begun to explore the application and performance of the Mamba model in various visual tasks. Representative works, such as the visual state space model [39] and vision Mamba [53], have demonstrated Mamba’s superiority for image processing, particularly in obtaining fewer parameters and faster speed. However, most Mamba-based models only conduct the scanning process along the spatial dimension, lacking cross-channel interaction, despite the importance of spectral information for HSI reconstruction. Recently, some Mamba-based methods have made progress in developing an SSM mechanism along the channel (spectral) dimension for certain visual tasks [43, 22], achieving better results compared to scanning only along the spatial dimension. However, this scanning method does not consider the inherently local features of HSIs in both spatial and spectral dimensions. Additionally, previous reconstruction algorithms have achieved better results in experiments by continuously improving the backbone and framework. However, it is often overlooked that achieving superior performance in the simulation experiment is not entirely linearly correlated with achieving better visual effects in the real experiment. Simulation experiments primarily address the losses caused by the reconstruction compression process, whereas real experiments introduce noise that simulates real-world conditions, thus requiring the model to possess stronger generalization capabilities.
However, the degradation caused by insufficient generalization capability is, in fact, not easy to notice. As a result, most current work has not recognized this issue. This oversight is mainly due to the limited test data from real-world experiments, which makes certain issues less likely to be noticed. In most scenes, an improvement in quantitative results in simulation experiments is indeed associated with better performance in real scenes, which has led to some confusion. In this paper, we provide a detailed explanation of this phenomenon, making the problem more apparent. Notably, certain optimization strategies commonly used in previous simulation experiments can even deteriorate specific features in real-world scenes. We assume this phenomenon arises due to the model’s overemphasizing of compression loss learning, thereby neglecting other noise (like shot noise) and the semantic information inherent in images. Experimental results in this paper confirm this point.
@@
In this paper, we propose CS-Mamba, a spatial-spectral SSM with our cross-scanning method and local enhancement, to improve HSI reconstruction in SCI. Our CS-Mamba aims to strike a balance between global and local information utilization in HSIs. CS-Mamba comprises two main components: spatial SSM and spectral-spatial cross SSM. The spatial SSM employs a combination of global and local scanning methods, enhancing row-by-row scanning with both global and local modeling. Meanwhile, the spectral-spatial cross SSM incorporates an innovative across-scanning mechanism, leveraging spatial-spectral local cubes to enhance the module’s ability to capture local details. Besides, we introduce a masked training strategy in real scenarios. This strategy forces the model to learn the inherent features of images, enhancing the generalization capabilities of previous methods and thereby achieving better results in real scenes. Overall, our contributions are summarized as follows:
@@
  • We propose a cross-scanning method that integrates local features from both spatial and spectral dimensions within a single scan. Additionally, we introduce a local scanning approach within the spatial SSM to further enhance local feature extraction.
  • We introduce a masked training strategy to promote the robustness of models in real scenarios.
Figure 2: Schematic of the CASSI system: A 3D HSI cube is modulated by a mask, sheared by a disperser, and transformed into a 2D-coded measurement.
@@ 2 Preliminaries and Related Works
CASSI System. An imaging process CASSI system is illustrated in Fig. 2. The 3D HSI data cube is denoted as $\mathbf{F}\in\mathbb{R}^{H\times W\times N_{\lambda}}$ where H, W, and $N_{\lambda}$ represent the height, width, and number of wavelengths of the 3D HSI, respectively. A physical mask $\mathbf{M}\in\mathbb{R}^{H\times W}$ is used to modulate the 3D HSI $\mathbf{F}$ as: $\mathbf{F^{{}^{\prime}}}(:,:,n_{\lambda})=\mathbf{F^{{}^{\prime}}}(:,:,n_{\lambda})\odot\mathbf{M},$ where $\odot$ represents the element-wise product and $\mathbf{F^{{}^{\prime}}}\in\mathbb{R}^{H\times W\times N_{\lambda}}$ is the modulated 3D HSIs. After the modulated cube passes the disperser, $\mathbf{F^{{}^{\prime}}}$ is tilted and considered to be sheared along the y-axis. We use $\mathbf{F^{{}^{\prime\prime}}}\in\mathbb{R}^{H\times(W+d(N_{\lambda}-1))\times N_{\lambda}}$ to denote the tilted cube and assume $n_{\lambda}$ to be the reference wavelength. That is, image $\mathbf{F^{{}^{\prime}}}(:,:,n_{\lambda_{c}})$ is not sheared along the y-axis. Then we have: $\mathbf{F^{{}^{\prime\prime}}}(u,v,n_{\lambda})=\mathbf{F^{{}^{\prime}}}(x,y+d(\lambda_{n}-\lambda_{c}),n_{\lambda})$, where $(u,v)$ indicates the coordinate system on the detector plane, and $\lambda_{n}$ is the wavelength of channel $n_{\lambda}$. Here, $d(\lambda_{n}-\lambda_{c})$ signifies the spatial shifting for band $n_{\lambda}$. Finally, the measurement $\mathbf{Y}\in\mathbb{R}^{H\times[W+d(N_{\lambda}-1)]}$ can be obtained by
@@
State Space Models. Recent advancements in state space models (SSMs) have spurred the development of various innovative methods and architectures. Notably, the structured state space model (S4) [18] has emerged as a pioneering approach due to its effectiveness in capturing long-range dependencies. S4 maps a 1D sequence $x(t)\in\mathbb{R}\rightarrow y(t)\in\mathbb{R}$ by an implicit latent state $h(t)\in\mathbb{R}^{N}$ and decodes it into an output sequence, which can be formulated as:
@@
where $N$ is the state size, ${\rm\textbf{A}}\in\mathbb{R}^{N\times N}$, ${\rm\textbf{B}}\in\mathbb{R}^{N\times 1}$, ${\rm\textbf{C}}\in\mathbb{R}^{1\times N}$, and ${\rm\textbf{D}}\in\mathbb{R}$. After that, a discretization process is adopted to integrate Eq. (2) into the deep learning algorithm, using the zero-order hold (ZOH) rule defined as,
@@
Mamba in Vision. Selective structured state space models (S6), or Mamba [17], use a selective mechanism for data-dependent dynamics, enhancing language and visual task performance. Vision Mamba (Vim) [53] addresses Mamba’s unidirectional modeling by adding bidirectional SSMs for global context and position embeddings for visual recognition. VMamba [28] bridges 1-D and 2-D scanning via the Cross-Scan Module (CSM). CU-Mamba [11] adds Channel SSM to capture channel correlations, while S2Mamba [43] introduces spectral scanning for semantic retrieval in HSI data. However, these methods overlook local similarity in both spatial and spectral dimensions of HSIs.
Figure 3: The flowchart of the proposed CS-Mamba. (a) Parameter Estimation Network. This module is used for capturing parameter $\alpha$ and noise level $\beta$ to guide the reconstruction process. (b) The U-shaped denoiser network and DUN framework. (c) The Spatial-Spectral SSM Block adopted in (b). This module mainly consists of a spatial SSM and a spectral SSM, depicted in the sub-flowchart (d) and (e), respectively. (f) The specific method of LE-SSM in the Spatial SSM Module. Scanning directions are divided into global and local branches. (g) The illustration of our Cross-Scan Mechanism. The scanning process is demonstrated by labeling the order of each data point.
3 The Proposed CS-Mamba Method
3.1 Architecture of CS-Mamba
The architecture of our CS-Mamba is illustrated in Fig. 3. CS-Mamba leverages a deep unfolding framework guided by mask degradation patterns. Using MAP theory, we model the HSI reconstruction process as follows:
$\mathbf{\hat{x}}=\text{arg}~{}\underset{\mathbf{x}}{\text{min}}~{}~{}\frac{1}{2}||\mathbf{y}-\mathbf{\Phi}\mathbf{x}||^{2}+\tau R(\mathbf{x}),$ (4)
where $||\mathbf{y}-\mathbf{\Phi}\mathbf{x}||^{2}$ is the data fidelity term, $R(\mathbf{x})$ represents the image prior, and $\tau$ controls the balance between the data fidelity and prior terms.
HQS introduces an auxiliary variable $z$, resulting in a constrained optimization problem given by
$\hat{x}=\arg\min_{x}\frac{1}{2}\|y-\Phi x\|^{2}+\lambda R(z)\quad s.t.\quad z=x.$ (5)
Eq. (5) is solved by minimizing
$\mathcal{L}_{\mu}(x,z)=\frac{1}{2}\|y-\Phi x\|^{2}+\lambda R(z)+\frac{\mu}{2}\|z-x\|^{2},$ (6)
where $\mu$ is a penalty parameter. Subsequently, Eq. (6) can be addressed by iteratively solving a data subproblem (x-subproblem) and a prior subproblem (z-subproblem):
${x_{k}=\arg\min_{x}\|y-\Phi x\|^{2}+\mu\|x-z_{k-1}\|^{2},}$ (7a)
${z_{k}=\arg\min_{z}\frac{1}{2(\sqrt{\lambda/\mu})^{2}}}\|z-x_{k}\|^{2}+R(z).$ (7b)
The data subproblem, Eq. (7a), usually has a fast closed-form solution as
$x_{k}=(\Phi^{\top}\Phi+\mu I)^{-1}(\Phi^{\top}y+\mu z_{k-1}),$ (8)
where $I$ is an identity matrix, the matrix inversion in Eq. (8) can be written as
$(\Phi^{\top}\Phi+\mu I)^{-1}=\mu^{-1}I-\mu^{-1}\Phi^{\top}(I+\Phi\mu^{-1}\Phi^{\top})^{-1}\Phi\mu^{-1}.$ (9)
In CASSI systems, $\Phi\in\mathbb{R}^{HW^{\prime}\times HW^{\prime}N_{\lambda}}$ is a block diagonal matrix, thus, $\Phi\Phi^{\top}$ is a diagonal matrix that can be defined as
$\Phi\Phi^{\top}\stackrel{{\scriptstyle def}}{{=}}\text{diag}\{\phi_{1},...,\phi_{i},...,\phi_{HW^{\prime}}\}.$ (10)
By plugging $\Phi\Phi^{\top}$ into $(I+\Phi\mu^{-1}\Phi^{\top})^{-1}$, we obtain:
$(I+\Phi\mu^{-1}\Phi^{\top})^{-1}=\text{diag}\{\frac{\mu}{\mu+\phi_{1}},...,\frac{\mu}{\mu+\phi_{i}},...,\frac{\mu}{\mu+\phi_{HW^{\prime}}}\}.$ (11)
By plugging Eq. (9), Eq. (10) and Eq. (11) into Eq. (8) and simplifying the formula, we have
$x_{k}=z_{k-1}+\Phi^{\top}[(y-\Phi z_{k-1})./(\mu+diag(\Phi\Phi^{\top}))],$ (12)
where $diag$ represents the extraction of elements from its diagonal to form a vector.
The prior subproblem Eq. (7b), from a Bayesian perspective, corresponds to Gaussian denoising on $x_{k}$ with noise level $\sqrt{\lambda/\mu}$. To address this, Eq. (7b) could be rewritten as,
$z_{k}=Denoiser(x_{k},\sqrt{\lambda/\mu}).$ (13)
DUNs iteratively solve the data subproblem (Eq. 12) and the prior subproblem (Eq. 13) to address ill-posed inverse problems. The data subproblem depends on the degradation matrix, which can be estimated using either a sensing matrix [3, 2] or a DNN [5]. However, sensing matrices may not accurately reflect real-world degradations, and directly modeling them is challenging. The prior subproblem is treated as a denoising task, but the noise level is unknown in reconstruction, affecting performance. Following [14], we iteratively update the degradation matrix and parameter estimates, learning constant noise and regularization parameters towards each stage due to computational cost, as detailed in the supplementary material.